Suppose you have decided to start a small business selling snacks from vending machines. You have secured a location for one candy vending machine in a local bookstore. Rental for the space will cost $200 per month.

Vending machines can be purchased at wholesale clubs such as Sam’s Club and Costco. You can also purchase the snacks to stock the machines in bulk there.

1. Either visit a local warehouse club or review its website to determine the initial cost to purchase a snack vending machine.

Vending machine from Sam’s Club: $4298.00

2. Assume you are initially going to have only one type of snack bar in your machine. What type of snack bar will you choose? If you purchase the bars in bulk, what is your cost per bar?

Snack Bar: Reese’s  

Reese’s in bulk: 36 pack for $28.92

Cost per bar: $0.80 (28.92/36)

3. How much will you charge for each bar sold?

$1.50  

4. What is your contribution margin per bar? How many bars must you sell to cover the cost of the vending machine?

5. When you have covered the initial investment, what will the monthly break-even point be in number of bars and in sales dollars?

6. Repeat requirements 2 through 5 assuming you decided to have a drink vending machine instead of a snack machine. Remember to find the price for a beverage vending machine. You may assume rental for the vending machine space is $200 per month regardless of its type.

7. Assume your machine can accommodate more than one product, for example, a snack machine that can dispense both chips and candy bars or a drink machine that offers both soda and water. Repeat requirements 2 through 5 assuming you have decided to offer two products. The products should have different prices, variable costs, and contribution margins. Perform the analyses for three levels of sales mix: 50/50, 30/70, and 70/30.

PLEASE ANSWER QUESTIONS 4-7 THANKS!
Answer #1

.4.Contribution Margin=(Sales Price) –(Variable Expenses)

Sales Price for each bar=$1.50

Variable expense=Cost per bar=$0.80

Contribution margin per bar=$1.50-$0.80=$0.70

Number of bars required to be sold to cover the cost of vending machine=4298/0.07=61,400

5.Monthly Rent =$200

Monthly Breakeven point in units=(Fixed Costs)/(Unit Contribution Margin)=200/0.07=2,857.143

Monthly Breakeven sales in units =2858(Rounded to whole units)

Monthly Breakeven Sales dollars=2858*1.50=$4,287